Demand in a market is given by Q=3p^-7.6 where p is the market price. What is the elasticity of demand? Include the negative sign if necessary.
1
Expert's answer
2019-03-25T07:32:27-0400

Q=3p^-7.6.

The elasticity of demand is:

"Ed=\frac{dQ}{dP}\times \frac{P}{Q} = 3\times (-7.6)\times p^{-8.6}\times \frac{p}{3 p^{-7.6}} = -7.6"